OpenStudy (agentnao):

The sum of n root 2 from 1 to infinity. I need desperately help with series! Please tell me you know how to find whether this is convergent or divergent?

OpenStudy (agentnao):

\[\sqrt[n]{2}\]

OpenStudy (anonymous):

the limit does not exist or it is infinite, then we say that the improper integral is divergent.

OpenStudy (agentnao):

Why does the limit not exist? As n approaches infinity, doesn't the value get closer to zero?

OpenStudy (anonymous):

Yes, but it will never get to zero. That's what makes it infinite

OpenStudy (agentnao):

But a limit doesn't ever actually REACH something, does it? It just gets reallllllly close.

OpenStudy (anonymous):

Exactly. Even though you are GETTING closer, you will NEVER reach zero

OpenStudy (agentnao):

Okay, wait. Let me give you another question then. \[\sum_{n=1}^{\infty} \frac{ 1 + 2^{n} }{3^{n}}\]

OpenStudy (agentnao):

How do I find if this is divergent or convergent? L'hospital's doesn't work, dividing by the highest power doesn't work, can't take the limit. So what else do I do?

OpenStudy (agentnao):

I'm sorry I'm struggling so much with this, but I just don't understand. I know one way to test for divergence is the divergence test. Meaning, if the limit of the series DNE or does not equal 0, then the series is divergent. However, a series CAN add up to a number other than 0 and still be convergent. What am I missing here?

OpenStudy (anonymous):

If anything is divergent, it just means it will never touch zero The sum of n root 2 from 1 to "infinity" In your statement infinity tells it all

OpenStudy (agentnao):

But all of my series go to infinity and some of them ARE convergent.

OpenStudy (agentnao):

For example: \[\sum_{n=1}^{\infty} \frac{ 1 }{ \sqrt{2}^{n}}\] is convergent.

OpenStudy (accessdenied):

In the first example, perhaps it would be useful to compare it to 1^(1/n). In your second example, \( \displaystyle \sum_{n=1}^{\infty} \dfrac{1 + 2^n}{3^n} \) If you understand that 1^n = 1, you might notice the opportunity for two geometric series here.

OpenStudy (accessdenied):

Where \( \sqrt[n]{2} = 2^{1/n} \) by definition. (If that doesn't make sense, think \( \left(2^{1/n}\right)^n = \left(\sqrt[n]{2}\right)^{n} \)... they both cancel the operation, right?)